# EffoVPR: Effective Foundation Model Utilization for Visual Place Recognition

### v1

@@ Appendix A Datasets
We evaluate EffoVPR performance across a large number of datasets, to underscore its top-performance in variable scenarios and cities. Following prior work, we have used [10] open-source code for downloading and organizing datasets, to ensure maximum reproducibility. In the following we shortly describe each of the datasets.
@@ A.1 Datasets Summary
AmsterTime[48] consists of $1,231$ image pairs from Amsterdam, Holland, exhibiting long-term changes. The queries are historical grayscale images, where for each query there is a reference of a modern-day photo which represents the same place. The pairs curated by human experts, and provide multiple challenges over different viewpoints and cameras, color vs grayscale and long-term changes.
Eynsham [15] is a collection of a car street-view camera, capturing photos around the same route of Oxford countryside twice. The grayscale images are divided to $23,935$ queries and $23,935$ gallery.
Mapillary Street-Level Sequences (MSLS) [47] is image and sequence-based VPR dataset. The dataset consists of more than 1.6M geo-tagged images collected during over seven years from 30 cities, in urban, suburban, and natural environments. There are 3 non-overlap subsets - a training set, validation (MSLS-val), and withheld test (MSLS-challenge). MSLS-val and MSLS-challenge provide various challenges, including viewpoint variations, long-term changes, and illumination and seasonal changes.
Nordland [41] was collected by a mounted camera on the top of a riding train in the Norwegian countryside, presenting rural and natural scenes. The data collected over the same route across four seasons, providing seasonal and illumination variability. Following [10, 41] we use the post-processed versions of winter as queries and summer as database, determining correct localization by retrieval of an image that is in less than 10 frames away. This dataset consists of $27,592$ query images and $27,592$ gallery images
Pittsburgh30k [4] is collected from Google Street View 360° panoramas of downtown Pittsburgh, split into multiple images. Ensuring queries and gallery were taken in different years, it provides 3 splits - a training set, validation and test. Pitts30k-test consists of 10k gallery images and 6816 queries. Pitts250k consists of $8280$ queries including these of Pitts30k, and its gallery size is $83,952$.
San Francisco Landmark (SF-R) [13] is a dataset from downtown San Francisco, which provides viewpoint variations. It presents a collection of $598$ of smartphone camera queries and gallery of $1,046,587$ images.
San Francisco eXtra Large (SF-XL) [8, 6] is an enormous dataset covering the whole city of San Francisco. it consists of a training set, which includes also raw 360° panoramas, a small validation set of $7,983$ queries and $8,015$ gallery images, and a test gallery of $2,805,840$ images.
@@
SF-XL-v1[8] consists of $1,000$ queries curated from Flickr, and provides viewpoint and camera variations, illumination changes and even some occlusions.
SF-XL-v2[8] is the queries of San Francisco Landmark (SF-R).
SF-XL-Night[6] is a collection of $466$ Flickr images of night scenes from San-Francisco. It provides viewpoint variations and very-challenging illumination changes.
SF-XL-Occlusion[6] is a collection of $76$ Flick images from the city of San Francisco, which suffers from severe occlusions, mostly by vehicles and crowd.
SPED[14] is a collection of surveillance cameras images consists of $607$ pairs of queries and gallery, captured accros time. It provides challenging viewpoint with seasonal and illumination changes.
St Lucia[36] is a collection of a nine videos of car-mounted camera from the St Lucia suburb of Brisbane. Following [10] open-source code, we select the first and last videos as queries and database, and sample one frame every 5 meters of driving. The gallery consists of $1,549$ images and there are $1464$ query images.
SVOX[12] is a dataset which presents multiple weather conditions VPR challenge. It consists of $17,166$ gallery images, of the city of Oxford. The queries were collected from the Oxford RobotCar dataset [35], providing multiple weather conditions queries sets, such as night (823 queries), overcast (872 queries), rainy (937 queries), snowy (870 queries) and sunny (854 queries).
Tokyo 24/7 [43] is a dataset from downtown Tokyo, which provides viewpoint changes and challenging illumination variations. It consists of a gallery of $75,984$ images, and a collection of $315$ smartphone camera queries from $185$ places. Each place is portrayed by three photos - one taken during the day, one at sunset and one at night.
@@ A.2 Train Dataset
Following VPR classification methods, Eigneplaces and CosPlace, we train on SF-XL[8] while other studies[1, 22, 33, 2] train on GSV-Cities[1] or combinations of Pittsburgh30k[4] and MSLS[47, 4, 34, 49], including a large mixture of different cities around the world (introducing higher variability). Note that similar to EigenPlace, our approach is designed for training on panoramas with heading information, and requires slicing them for lateral and frontal views, which cannot be applied other training datasets.
Appendix B Ablation Study
Here we conduct extensive experiments on two different datasets to ablate over several key-components of our EffoVPR method.
@@ Appendix B Ablation Study
Number of Candidates to Re-rank: The second re-ranking stage is applied to the top-K candidates retrieved during the global stage. Although common choice in literature is $K=100$ (e.g. [34, 49, 10]), we explore different choices of K, as detailed in Table S4. We achieve SoTA results even with $K=5$. It is important to note that in some cases, an increase in K can introduce a greater number of “distractor” candidates, potentially leading to a decrease in performance. However, EffoVPR SoTA performance is consistent for all tested K’s.
@@
Choice of Trainable Layers: Table S5 presents a few different sets of trainable layers in our backbone model. We find the vanilla fine-tuning of the entire model, end-to-end, that includes all layers, drastically harms the performance of EffoVPR. We attribute this decline to the fact that the DINOv2 backbone was trained on significantly larger datasets compared to those typically used in VPR. Subsequently, we establish that training only the last five layers represents a “sweet-spot”, yielding peak performance. Both increasing or decreasing the number of trainable layers from this configuration leads to lower results.
Table S 5: Ablation study on the choice of trainable last layers. Results (in %) are the R@1 of the 1st stage. Note that the 0 column represent a zero-shot performance of DinoV2.
Dataset 0 1 2 3 4 5 6 all layers
MSLS-val 47.7 89.5 88.2 89.7 89.5 90.9 89.7 86.1
Tokyo-24/7 62.2 96.8 96.5 95.9 96.2 97.5 96.8 94.0
@@
Fig. S 1: Recall@1 performance of EffoVPR-G global feature versus feature dimensionality for more datasets.
@@ D.1 Additional Zero-Shot Visualizations
Fig. S 2: Additional EffoVPR-ZS zero-shot visualizations.
@@ D.2 Additional Re-ranking Visualizations
Fig. S 3: EffoVPR-R top-1 matching visualizations. For each pair matching, the left image is the query and the right is the top-1 result.
@@ Appendix E Implementation details
We use ViT-L/14 as the backbone, initialized with pre-trained weights of DINOv2 with registers[37, 16]. We only train the last five layers of the backbone, which appeared to be most beneficial. We employ EigenPlaces’s[11] group and class partitioning with its default hyper-parameters, and both lateral and frontal views, on the publicly available SF-XL street-view panoramas dataset [8]. We set an AdamW optimizer to the backbone, and an Adam optimizer to the classification heads, both with a constant learning rate of $1\times 10^{-5}$. We train EffoVPR with a batch size of 16, for 25 epochs, on a single NVIDIA-A100 node. We otherwise follow EigenPlaces training recipe. We choose the best epoch by SF-XL validation set, measuring Recall@1 global ranking performance. Given that ViT is independent of the image input size (provided it can be segmented into $14\times 14$ patches), we evaluated using images sized $504\times 504$, but trained on 224 × 224 images to expedite training. For benchmarking EffoVPR-G of the global feature, we report nearest-neighbors performance on normalized output class token. In the re-ranking stage we extract the $V$ self-attention facet from layer $n-1$ (with $n$ being the output layer), measure cosine similarity, we filter the features by class attention map with a threshold $\mathcal{T}_{1}=0.05$, and count only mutual nearest-neighbors with a score above the threshold $\mathcal{T}_{2}=0.65$.
@@ E.1.1 Zero-shot
In evaluating AnyLoc[27], we tackle the significant memory requirements of its VLAD pooling algorithm by implementing an online clustering scheme. We observed that their recommendation for layer 31 outperformed layer 23. In our zero-shot evaluation, we assess EffoVPR-ZS method by extracting the $V$ features from layer $n-2$ to re-rank the top-100 candidates retrieved from the first-stage global [CLS] feature. In this framework, our performance is constrained by the first-stage Recall@100, achieving rates of $99.2\%$, $96.8\%$, $81.5\%$, and $78.1\%$ on Pitts30k, Tokyo24/7, MSLS-Val, and Nordland, respectively.
@@ E.1.2 Benchmarking
Generally, for consistent benchmarking, we adhere to [10]. In addition, we report the results of other methods in accordance with the evaluation choices of SelaVPR[34] and CricaVPR[33], including the specific versions of trained models utilized. For the recent state-of-the-art methods SelaVPR, CricaVPR, and SALAD, we provide results from the original publications whenever available. When such results are not directly available, we utilize their code and published weights. Specifically for SelaVPR, which has two sets of weights (trained on Pitts30k and MSLS), we report the best-performing for each dataset.
@@ Appendix F Additional Quantitative Results
Table S 6: Comparison to SoTA on more datasets

### v2

@@ Appendix A Datasets
We evaluate EffoVPR performance across a large number of datasets, to underscore its top-performance in variable scenarios and cities. Following prior work, we have used (Berton et al., 2022c) open-source code for downloading and organizing datasets, to ensure maximum reproducibility. In the following we shortly describe each of the datasets.
@@ A.1 Datasets Summary
AmsterTime (Yildiz et al., 2022) consists of $1,231$ image pairs from Amsterdam, Holland, exhibiting long-term changes. The queries are historical grayscale images, where for each query there is a reference of a modern-day photo which represents the same place. The pairs curated by human experts, and provide multiple challenges over different viewpoints and cameras, color vs grayscale and long-term changes.
Eynsham (Cummins & Newman, 2009) is a collection of a car street-view camera, capturing photos around the same route of Oxford countryside twice. The grayscale images are divided to $23,935$ queries and $23,935$ gallery.
Mapillary Street-Level Sequences (MSLS) (Warburg et al., 2020) is image and sequence-based VPR dataset. The dataset consists of more than 1.6M geo-tagged images collected during over seven years from 30 cities, in urban, suburban, and natural environments. There are 3 non-overlap subsets - a training set, validation (MSLS-val), and withheld test (MSLS-challenge). MSLS-val and MSLS-challenge provide various challenges, including viewpoint variations, long-term changes, and illumination and seasonal changes.
Nordland (Sünderhauf et al., 2013) was collected by a mounted camera on the top of a riding train in the Norwegian countryside, presenting rural and natural scenes. The data collected over the same route across four seasons, providing seasonal and illumination variability. Following (Berton et al., 2022c; Sünderhauf et al., 2013) we use the post-processed versions of winter as queries and summer as database, determining correct localization by retrieval of an image that is in less than 10 frames away. This dataset consists of $27,592$ query images and $27,592$ gallery images
Pittsburgh30k (Arandjelovic et al., 2016) is collected from Google Street View 360° panoramas of downtown Pittsburgh, split into multiple images. Ensuring queries and gallery were taken in different years, it provides 3 splits - a training set, validation and test. Pitts30k-test consists of 10k gallery images and 6816 queries. Pitts250k consists of $8280$ queries including these of Pitts30k, and its gallery size is $83,952$.
San Francisco Landmark (SF-R) (Chen et al., 2011) is a dataset from downtown San Francisco, which provides viewpoint variations. It presents a collection of $598$ of smartphone camera queries and gallery of $1,046,587$ images.
San Francisco eXtra Large (SF-XL) (Berton et al., 2022a; Barbarani et al., 2023) is an enormous dataset covering the whole city of San Francisco. it consists of a training set, which includes also raw 360° panoramas, a small validation set of $7,983$ queries and $8,015$ gallery images, and a test gallery of $2,805,840$ images.
@@
SF-XL-v1 (Berton et al., 2022a) consists of $1,000$ queries curated from Flickr, and provides viewpoint and camera variations, illumination changes and even some occlusions.
SF-XL-v2 (Berton et al., 2022a) is the queries of San Francisco Landmark (SF-R).
SF-XL-Night (Barbarani et al., 2023) is a collection of $466$ Flickr images of night scenes from San-Francisco. It provides viewpoint variations and very-challenging illumination changes.
SF-XL-Occlusion (Barbarani et al., 2023) is a collection of $76$ Flick images from the city of San Francisco, which suffers from severe occlusions, mostly by vehicles and crowd.
SPED (Chen et al., 2018) is a collection of surveillance cameras images consists of $607$ pairs of queries and gallery, captured accros time. It provides challenging viewpoint with seasonal and illumination changes.
St Lucia (Milford & Wyeth, 2008) is a collection of a nine videos of car-mounted camera from the St Lucia suburb of Brisbane. Following (Berton et al., 2022c) open-source code, we select the first and last videos as queries and database, and sample one frame every 5 meters of driving. The gallery consists of $1,549$ images and there are $1464$ query images.
SVOX (Berton et al., 2021) is a dataset which presents multiple weather conditions VPR challenge. It consists of $17,166$ gallery images, of the city of Oxford. The queries were collected from the Oxford RobotCar dataset (Maddern et al., 2017), providing multiple weather conditions queries sets, such as night (823 queries), overcast (872 queries), rainy (937 queries), snowy (870 queries) and sunny (854 queries).
Tokyo 24/7 (Torii et al., 2015) is a dataset from downtown Tokyo, which provides viewpoint changes and challenging illumination variations. It consists of a gallery of $75,984$ images, and a collection of $315$ smartphone camera queries from $185$ places. Each place is portrayed by three photos - one taken during the day, one at sunset and one at night.
@@ A.2 Train Dataset
Following VPR classification methods, Eigneplaces and CosPlace, we train on SF-XL (Berton et al., 2022a) while other studies (Ali-bey et al., 2022; Izquierdo & Civera, 2024; Lu et al., 2024a; Ali-bey et al., 2023) train on GSV-Cities (Ali-bey et al., 2022) or combinations of Pittsburgh30k (Arandjelovic et al., 2016) and MSLS (Warburg et al., 2020; Arandjelovic et al., 2016; Lu et al., 2024b; Zhu et al., 2023), including a large mixture of different cities around the world (introducing higher variability). Note that similar to EigenPlaces, our approach is designed for training on panoramas with heading information, and requires slicing them for lateral and frontal views, which cannot be applied other training datasets.
Appendix B Ablation Study (More)
We conduct extensive experiments on two different datasets to ablate over key-components of our EffoVPR method.
@@ Appendix B Ablation Study (More)
Number of Candidates to Re-rank: The second re-ranking stage is applied to the top-K candidates retrieved during the global stage. Although common choice in literature is $K=100$ (e.g. (Lu et al., 2024b; Zhu et al., 2023; Berton et al., 2022c)), we explore different choices of K, as detailed in Table S4. We achieve SoTA results even with $K=5$. It is important to note that in some cases, an increase in K can introduce a greater number of “distractor” candidates, potentially leading to a decrease in performance. However, EffoVPR SoTA performance is consistent for all tested K’s.
@@
Choice of Trainable Layers: Figure S1 presents a few different sets of trainable layers in our backbone model. We find the vanilla fine-tuning of the entire model, end-to-end, that includes all layers, drastically harms the performance of EffoVPR. We attribute this decline to the fact that the DINOv2 backbone was trained on significantly larger datasets compared to those typically used in VPR. Subsequently, we establish that training only the last five layers represents a “sweet-spot”, yielding peak performance. Both increasing or decreasing the number of trainable layers from this configuration leads to lower results.
Fig. S 1: Ablation on the number of trainable layers
$T_{1}$ and $T_{2}$ thresholds. $T_{1}$ parameter shown in Fig.2 sets a threshold for selecting strong keypoints, while $T_{2}$ sets a threshold for identifying strong matching pairs. In practice, keypoints with scores ($s_{i}$ in Fig. 2) below $T_{1}$ are filtered out, and matching pairs with scores below $T_{2}$ are discarded. We set $T_{1}$ and $T_{2}$ values according to a validation set, and ablate the sensitivity of the values in Tables SLABEL:tab:T1_ablation, SLABEL:tab:T2_ablation.
Table S 5: Ablation study on the value of $T_{1}$. Results (in %) are the R@1.
$T_{1}$ 0.00 0.05 0.10
Tokyo24/7 97.8 98.7 98.4
MSLS-val 89.1 92.8 92.4
Table S 6: Ablation study on the value of $T_{2}$. Results (in %) are the R@1.
$T_{2}$ 0.50 0.65 0.80
Tokyo24/7 97.5 98.7 98.4
MSLS-val 91.8 92.8 91.5
Image Resolution. Different methods employed varying image resolutions. For instance, SALAD (Izquierdo & Civera, 2024) used a resolution of 322px. During inference, we used images with a resolution of 504px to enable more patches (of $14\times 14$), particularly for re-ranking that uses local keypoint detection and matching.
The difference in inference time is in the feed-forward path of a larger input through the transformer. Our observations show that the feed-forward time for 224px image is $13.82$ milliseconds and for $504$px image is $28.8$ milliseconds on average (with a single A100 node). Note that the local keypoints obtain the same size of descriptors so the matching time and memory consumption are of the same order for larger input images.
Table S7 presents an ablation study on performance using intermediate resolutions of $224$px and $322$px on the Tokyo24/7 and MSLS-val datasets. EffoVPR continues to demonstrate state-of-the-art results (Recall@1), while showing slight improvement with $504$px.
Table S 7: Ablation study on image resolution. Results (in %) are the R@1
Res Tokyo24/7 MSLS-Val
$224$px 98.4 91.1
$322$px 98.4 92.8
$504$px 98.7 92.8
@@
Fig. S 2: Recall@1 performance of EffoVPR-G global feature versus feature dimensionality for more datasets.
@@ D.1 Additional Zero-Shot Visualizations
Fig. S 3: Additional EffoVPR-ZS zero-shot visualizations.
@@ D.2 Additional Re-ranking Visualizations
Fig. S 4: EffoVPR-R top-1 matching visualizations. For each pair matching, the left image is the query and the right is the top-1 result.
@@ Appendix E Implementation details
We use ViT-L/14 as the backbone, initialized with pre-trained weights of DINOv2 with registers (Oquab et al., 2023; Darcet et al., 2023). We only train the last five layers of the backbone, which appeared to be most beneficial. We employ EigenPlaces’s (Berton et al., 2023) group and class partitioning with its default hyper-parameters, and both lateral and frontal views, on the publicly available SF-XL street-view panoramas dataset (Berton et al., 2022a). We set an AdamW optimizer to the backbone, and an Adam optimizer to the classification heads, both with a constant learning rate of $1\times 10^{-5}$. We train EffoVPR with a batch size of 16, for 25 epochs, on a single NVIDIA-A100 node. We otherwise follow EigenPlaces training recipe. We choose the best epoch by SF-XL validation set, measuring Recall@1 global ranking performance. Given that ViT is independent of the image input size (provided it can be segmented into $14\times 14$ patches), we evaluated using images sized $504\times 504$, but trained on 224 × 224 images to expedite training. For benchmarking EffoVPR-G of the global feature, we report nearest-neighbors performance on normalized output class token. In the re-ranking stage we extract the $V$ self-attention facet from layer $n-1$ (with $n$ being the output layer), measure cosine similarity, we filter the features by class attention map with a threshold $\mathcal{T}_{1}=0.05$, and count only mutual nearest-neighbors with a score above the threshold $\mathcal{T}_{2}=0.65$.
@@ E.1.1 Zero-shot
In evaluating AnyLoc (Keetha et al., 2023), we tackle the significant memory requirements of its VLAD pooling algorithm by implementing an online clustering scheme. We observed that their recommendation for layer 31 outperformed layer 23. In our zero-shot evaluation, we assess EffoVPR-ZS method by extracting the $V$ features from layer $n-2$ to re-rank the top-100 candidates retrieved from the first-stage global [CLS] feature. In this framework, our performance is constrained by the first-stage Recall@100, achieving rates of $99.2\%$, $96.8\%$, $81.5\%$, and $78.1\%$ on Pitts30k, Tokyo24/7, MSLS-Val, and Nordland, respectively.
@@ E.1.2 Benchmarking
Generally, for consistent benchmarking, we adhere to (Berton et al., 2022c). In addition, we report the results of other methods in accordance with the evaluation choices of SelaVPR (Lu et al., 2024b) and CricaVPR (Lu et al., 2024a), including the specific versions of trained models utilized. For the recent state-of-the-art methods SelaVPR, CricaVPR and SALAD, we provide results from the original publications whenever available. When such results are not directly available, we utilize their code and published weights. Specifically for SelaVPR, which has two sets of weights (trained on Pitts30k and MSLS), we report the best-performing for each dataset. For BoQ, we evaluated by the model provided weights across all datasets, due to inconsistency in reported results in the paper, and project page and code.
@@ Appendix F Additional Quantitative Results
Table S 8: Comparison to SoTA on more datasets